OpenStudy (anonymous):

a recipe for a pie calls for 1/3 cup of blackberries and 2/6 cup of rasberries needed for the recipe ?

OpenStudy (anonymous):

question doesn't make sense...do you mean how many berries in total required? If yes, then: 1/3 + 2/6 find a common denominator = 2/6+2/6 =4/6 =2/3 cups